OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 8, -14, and 3 + 9i A. f(x) = x4 - 11x3 + 72x2 - 606x + 10,080 B. f(x) = x4 - 303x2 + 1212x - 10,080 C. f(x) = x4 - 11x3 - 72x2 + 606x - 10,080 D. f(x) = x4 - 58x2 + 1212x - 10,080

OpenStudy (cwrw238):

it wil have (x +14)(x - 8) as factors there will be another zero 3 - 9i

OpenStudy (cwrw238):

divide these factors into the polynomial

OpenStudy (anonymous):

im not actually sure how to do tht

OpenStudy (cwrw238):

right - i suspect its the last one D you can use synthetic division also but i don't like that (x +14)(x - 8) = x^2 + 6x - 112 x^2 -6x + 90 ------------------ x^2+ 6x-112}x^4- 0x^3 - 58x^2 + 1212x - 10080 x^4 +6x^3 -112x^2 ------------------ subtract -6x^3 + 54x^2 + 1212x -6x^3 - 36x^2 + 672x ------------------- subtract 90x^2 + 540x - 10080 90x^2 + 540x - 10080

OpenStudy (anonymous):

ohh i see okay, th

OpenStudy (anonymous):

thank you*

OpenStudy (cwrw238):

it was the correct one so D = (x+14)(x-8)(x^2 -6x + 90) the last part will have complex roots x = 6 +/- sqrt(36 - 4*90) / 2 which will give you 3+ 9i and 3 + 9 i

OpenStudy (anonymous):

i see, and totally get it now! thank you so much

OpenStudy (cwrw238):

yw